40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 15 02:32:01 2014, Jenne, Update, LSC, Started engaging the AO path, not getting all the way yet TFSR785_15-05-2014_011008.pdf
    Reply  Thu May 15 02:52:51 2014, Jenne, Update, LSC, Started engaging the AO path, not getting all the way yet 
Message ID: 9956     Entry time: Thu May 15 02:32:01 2014     Reply to this: 9957
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Started engaging the AO path, not getting all the way yet 

I tried many times this evening to engage the AO path, with limited success.

Q's new scripts worked really well, and so I have some transfer functions!  To take these measurements, in ...../scripts/general/netgpib, I am running ./TFSR785 TFSR785_CARMloop_May2014.yml, where the file name is the name of my parameter file.  The data, and the saved pdfs, are in /users/jenne/PRFPMI/CARM_loop_measurements/2014May14/ .  For these measurements, the SR785 is hooked up to the "A" set of excitation and test points on the CM board.

All of these traces were taken while the IFO was PRFPMI, with PRCL and MICH on REFL33, DARM on ALS diff, and CARM on InvSqrtTrans.  carm_cm_up.sh is up to date, through the echo "REFL_I should now be zero" (~line 111 in the script).  All you need to do is set the beatnotes, and then run the script.  Follow instructions in the prompt (such as "press enter to confirm PRMI is locked"). 

TFSR785_15-05-2014_011008.pdf

Here are my notes for the various times:

23:01:44 - MC IN2 = 0dB, CARM gain = 5.0

23:13:45 - MC IN2 = 10dB, CARM gain = 5

23:26:10 - MC IN2 = 6dB, CARM gain = 10 (after Q suggested increasing overall gain, rather than just AO path)

00:13:07 - MC IN2 = 6dB?, CARM = 6ish?  don't remember exactly.

00:45:00ish, Realigned IFO using IR with arms.

01:03:17 - MC In2 = 0dB, CARM gain = 5

01:07:42 - MC IN2 = 8dB, CARM gain = 6.295  (AO went up to 6dB, then +1dB steps to both simultaneously using  ezcastep C1:LSC-CARM_GAIN 1dB C1:IOO-MC_AO_GAIN 1)

01:08:57 - MC IN2 = 10dB, CARM gain = 7.92447

01:10:08 - MC IN2 = 12dB, CARM gain = 9.97631

lockloss when trying to add 1 more dB to both.

01:41:36 - MC IN2 = 12dB, CARM gain = 9.97

lockloss when just MC IN2 up by 1dB, left CARM gain alone.


Other notes:

The 60Hz noise in TRY is back.  Since I thought I remembered someone suggesting that it was leakage light from the exit sign, Koji went in and wrapped the end table in foil, however the lines are still present.

ELOG V3.1.3-